PRODUCTION OF A BEAM OF HIGHLY VIBRATIONALLY EXCITED CO USING PERTURBATIONS

\underline{N. BARTELS}, T. SCHÄFER, J. HÜHNERT, A. M. WODTKE, Georg August Universität Göttingen, Institut für Physikalische Chemie, Tammanstrasse 6, 37075 Göttingen, Germany; and R. W. FIELD, Massachusetts Institute of Technology, 77 Massachusetts Avenue, Cambridge, MA 02139, USA.

For many experimentalists (especially those, who are not spectroscopists), molecular pertubations are a curse, as they make assignments and analysis of spectral data more difficult. Nevertheless, they can also be a boon! In this talk we will show how a molecular beam of CO in high vibrational states (v = 17, 18) can be prepared by an optical pumping scheme that we call PUMP-PUMP-PERTURB and DUMP (P^{3}D). P^{3}D exploits the loaning, via spin-orbit perturbations, of the large oscillator strength of the 4th positive system, A^{1}Π ← X^{1}Σ^{+}, to the triplet manifold. This allows some nominally spin-forbidden transitions to be exploited in multistep optical pumping schemes. The ability to state-selectively prepare CO in high vibrational states opens up new opportunities for molecular beam scattering experiments.